OpenStudy (anonymous):

please help me with systems of equations! 2x-4y=12 3x-6y=21 (I need to find x and y)

OpenStudy (anonymous):

divided 2nd equation by 3 to get x-2y=10.5 x=10.5+2y then substitute by x=10.5+2y in 1st equation

OpenStudy (anonymous):

so 2(10.5+2)-4y=12? @esamalaa but how do I get x and y values?